Do you multiply or divide for this question, " Out of 18 times at bat, Emily got a hit 5/6 of the time and Julie got a hit 8/9 of the time. Who got more hits? How many more? " ?
1
Expert's answer
2019-10-25T11:31:27-0400

Total hit by Emily "=\frac{5}{6}\times 18=15" hits


Total hit by Julie "=\frac{8}{9}\times 18=16" hits

So,

No. Of hit by Juli are more than that of Emily by 16-15=1 hit.